Circular No. 5847

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 L. Wang, J. Y. Hu, A. G. Li, and H. B. Li, Beijing Astronomical Obser-

vatory; and Z. W. Li, Beijing Normal University, report: “Low-resolution spectrograms (∼ 0.53 nm/pixel; range 425-700 nm), taken on Aug. 8-9 with the 2.16-m telescope at the Xinglong station of Beijing Observatory, show that SN 1993J has now entered a nebular phase. The [O I] doublet at 630 and 636.4 nm is very strong and now overtakes the neighboring Hα and He I 667.8-nm complex in strength. [O I] at 557.7 nm is also a strong feature in the spectra. The [O I] 557.7-nm line is clearly blueshifted, its measured centroid velocity being ∼ –1775 km/s (FWHM ∼ 4300 km/s). The [O I] doublet (630.0, 636.4 nm) can be approximately fitted by the profile of the [O I] 557.7-nm line, assuming the [O I] 630.0- and 636.4-nm lines are of equal strength. The success of this profile fit means that the [O I] doublet is optically thick. The [O I] 630.0- and 636.4-nm lines are also clearly blueshifted, by about the same amount as measured in the [O I] 557.7-nm line. The double peak in Hα has now disappeared, while the overall width of the line remains the same. Many fine structures appeared in the Hα and [O I] lines. Typical FWHM of the feature is ∼ 750 km/s. The blueshift of the [O I] lines can be traced to even earlier dates, to as early as Apr. 28, ever since the [O I] line appeared. The fine structures clearly indicate that strong instabilities have occurred in the ejecta. The blueshift may be due to dust formation in the ejecta, as was observed by Danziger et al. in SN 1987A (IAUC 4746). However, it should be recalled that the blueshift in SN 1993J was observed at a much earlier date, and the amount of blueshift is much larger. It is not clear whether the ejecta can be cool enough for dust formation to occur.”

A. Clocchiatti, J. C. Wheeler, and D. A. Swartz report: “A fully reduced spectrogram (range 450-760 nm, dispersion 0.4 nm/pixel) of SN 1993J, obtained on Aug. 14.16 UT at the 2.1-m telescope of McDonald Observatory, shows that Hα and Hβ are still present in blueshifted absorption (8650 and 8800 km/s, respectively), superimposed on a very blue continuum. He I at 667.8 nm has disappeared and Hα emission nearly so. [O I] at 557.7 nm (centroid blueshifted by 1700 km/s) and [O I] at 630.0 and 636.5 nm (centroid blueshifted by 500 km/s, with a weaker, narrow component blueshifted by 2100 km/s) show fine structure in their profiles, confirming the Beijing report above.”

1993 August 18 Daniel W. E. Green